OpenStudy (anonymous):

What is the sum of a 22-term arithmetic sequence where the first term is 55 and the last term is -29?

OpenStudy (anonymous):

the formula for this is S = (n/2)[ a + l] in this case n = 22, a = 55 and l = -29

OpenStudy (anonymous):

so 286?

OpenStudy (anonymous):

11 * 26 = 286 yep

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

yw